# CrowdVLM-R1: Expanding R1 Ability to Vision Language Model for Crowd Counting using Fuzzy Group Relative Policy Reward

### v1

@@ IV-A Dataset
The dataset used in this article is derived from existing datasets for counting tasks. The training dataset includes aerial sheep [43], video game characters from GTAV_Head [44], street pedestrians [27], wheat heads from the Global Wheat Head Detection (GWHD) dataset [45] and cars from TRANCOS[46]. The testing dataset includes an additional dataset, manatee [47], for out-of-domain evaluation.
@@
Within our dataset, not all the images from original datasets were used. If original dataset has train and test sets, then the original train and test samples were picked randomly to use as part of our own train and test set, respectively. If the original data set was not split into train and test set, then the dataset was split into two parts first and then images from these two subsets were picked later into our train and test sets. For our dataset dataset, up to 1000 images and 100 samples were choose for each of the original datasets.
Table I provides a summary of the datasets used in our study, highlighting the diversity and complexity of the counting task. The dataset encompasses a wide range of scenarios, from relatively simple images containing only a few objects to highly complex scenes with thousands of densely packed, overlapping instances. The variation in object counts, as indicated by the minimum, maximum, and mean values, demonstrates the challenge of accurately estimating quantities across different contexts. Some datasets, such as the video game characters and street pedestrians, include images where objects are densely clustered, making separation and identification difficult. Additionally, environmental noise, including buildings, trees, and other background elements, further complicates detection in real-world scenes. The dataset complexity indicates the challenges of doing counting task for a single model.
@@ IV-B1 Models
  • Baseline Models: We evaluate several widely used benchmark models, including GPT-4o-2024-11-20 (GPT-4o), Claude Sonnet-3.5-2024-06-20 (Sonnet-3.5), LLaMA 2-90B (LLaMA-90B), Intern-VL-26B (In-VL2-26B), Qwen2.5-VL-3B (Q-3B), and Qwen2.5-VL-7B (Q-7B).
@@
  • GRPO Models: We conduct full-scale reinforcement learning (RL) training with the GRPO method on Qwen2.5-VL-3B (Q-3B-R1) and Qwen2.5-VL-7B (Q-7B-R1), employing the standard accuracy-based reward defined in Equ. 3.
  • CrowdVLM-R1 Models (Ours): We also perform full-scale RL training using our proposed FGRPR method on Qwen2.5-VL-3B (Q-3B-FR) and Qwen2.5-VL-7B (Q-7B-FR). Our fuzzy reward function, defined in Equ. 4.
@@ IV-B2 Experimental Environment
During training and testing for GRPO and FRGPR, we use 8 Nvidia L40S GPU (each 46GB memory). The training dataset consists of 3803 images and 5 classes, the experiment is set have max prompt length 1024, number of generations 8, per device training batch size 6, gradient accumulation steps 2, data types bf16, and epochs 2. The testing consists of 600 images and 6 classes (one class is out of training samples), we uses batch size 2 and data types bf16. It is worth noting that we use the same training and testing hyperparameters and prompt method across all the reinforcement learning models.
@@ IV-C Evaluation Metrics
To evaluate the performance of crowd counting task, we employ Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), which are standard metrics in the field. MAE measures the average absolute difference between predicted and actual crowd counts, providing an intuitive measure of accuracy:
@@
In order to have a comprehensive analysis, we also employee $R^{2}$, MAPE, Pearson’s correlation, and Spearman’s correlation. Their formal definitions can be found in Section VI-A.
@@ IV-D Results
The overall model performance across different domains is shown in Table II. Among the baseline models, LLaMA-90B demonstrates outstanding performance in four of the six domain datasets. Nevertheless, GPT-4o and Sonnet-3.5 achieve lower MAE and RMSE scores compared to LLaMA-90B. Specifically, GPT-4o exhibits comparable performance to LLaMA-90B on the out-of-domain dataset (manatee), displaying an MAE of approximately 5.7 and an RMSE around 10. On the other hand, Q-3B and Q-7B consistently show among the poorest metric results across all six datasets.
@@
However, after traditional supervised fine-tuning (SFT), both Q-3B-SFT and Q-7B-SFT improve significantly. Specifically, they outperform not only their base versions but also all baseline models, including state-of-the-art models such as GPT-4o and LLaMA-90B, across five in-domain datasets. Although Q-7B-SFT achieves a lower overall MAE compared to Q-3B-SFT, its performance is not consistently better across these five in-domain datasets. Furthermore, Q-7B-SFT presents a higher RMSE value, indicating that its prediction errors exhibit greater variance and instability.
@@
After fine-tuning with GRPO methods employing the traditional $0/1$ reward function (rewarding predictions with $1$ if $\frac{\hat{y}-y}{y}<0.5$, otherwise 00), the results indicate that although this approach has some effectiveness, the models Q-3B-R1 and Q-7B-R1 show worse performance than their corresponding directly fine-tuned counterparts, Q-3B-SFT and Q-7B-SFT. This outcome reveals a limitation of applying such a GRPO-based reward approach specifically for the crowd counting task.
@@
Models trained with our proposed FGRPR method, particularly Q-7B-FR, exhibit state-of-the-art performance across all five in-domain datasets. They continue to reduce the MAE and RMSE compared to the previously best-trained model, Q-7B-SFT, showing consistency in error reduction. For the 3B variant, Q-3B-FR achieves better performance than the previous SFT model (Q-3B-SFT) in two datasets, slightly worse performance in sheep counting, and equivalent performance in wheat head and traffic car counting datasets. Moreover, for the out-of-domain dataset, SFT and FGRPR shows similar performance that they improved compared to their based models that shows better performance than bigger models In-VL2-26B and Sonnet-3.5 but still can not achieve to the best industry models like GPT-4o and LLaMA-90B. Considering the small size of our base models(only 3B and 7B), these results collectively prove the effectiveness of our proposed FGRPR reward function.
This can be further confirmed by Table III that models are evaluated by the ranges. In domain datasets, while SFT and FGRPR methods have similar in relatively few objects counting, range [1-20] and [21-50], when object continue growing over 50, FGRPR shows dominant performance over SFT and GRPO which controlled by accuracy reward, Equ.3, of $r_{a}$. This shows the effectiveness of our reward function, Equ.7, the part of $\frac{\left|\hat{y}-y\right|}{y}$ which provides higher reward with better counting value.
@@
This effectiveness is further supported by Table III, where the models are evaluated based on different count ranges. In in-domain datasets, the SFT and FGRPR methods have similar performance with fewer objects, in ranges [1-20] and [21-50]. However, as object counts exceed 50, FGRPR shows superior performance compared to SFT and GRPO, which is guided by the accuracy reward, Equ. LABEL:equ:ra. This highlights the effectiveness of our proposed reward function, Equ. LABEL:equ:rp, particularly the component ($\frac{\left|\hat{y}-y\right|}{y}$), which offers higher rewards for more accurate counts.
@@ IV-E1 Model rewards
Figure 5: Training reward curves for different models show the same trend while for precision reward, bigger models dominate the smaller models within the same method over the training steps after smoothed. And their format reward lines do not have any difference after 100 steps.
In Fig. 5, all four models demonstrate an upward trend in both precision and format rewards, as shown in the first-row subfigures where a simple moving average with a window size of 20 is applied across 600 training steps. The 7B models start at higher points than the 3B models, indicating superior initial performance, consistent with prior studies that larger models typically outperform smaller ones within the same family. While 3B with FGRPR has similar start values with Q-7B-R1.
All four models display a sharp increase within the initial 100 training steps, followed by fluctuating yet steady improvement in precision rewards over the subsequent 500 steps, reaching their peak around step 600. Notably, the FGRPR method demonstrates a sharper and steadier growth trajectory compared to the traditional GRPO method. Furthermore, the continuous upward trend shown by the FGRPR model at the training endpoint suggests potential for further improvement with additional training data or a longer training period, whereas models trained via GRPO with traditional rewards plateau and exhibit little to no increase in reward values during the second half of training. Meanwhile, all four models have identical format reward curves, as they share the same format reward function.
@@ IV-E2 Concentrated results
Fig. 6 presents distribution statistics for ground truth over prediction results. The histogram and density plots show both distributions are highly right-skewed, with the majority of values concentrated below 250. However, the second row subfigures dramatically illustrate disparity between GT and predictions - GT values show numerous outliers spanning a much wider range, while predicted values display fewer outliers and compressed variability.
@@
Table IV reveals additional metrics for the prediction of the model. The $R^{2}$ value of 0.8795 indicates that the model explains nearly 88% of the variance. Strong correlation coefficients (Pearson = 0.9655, Spearman = 0.9497) confirming the model generally preserves rank ordering. And A MAPE of 23.89% suggests predictions are, on average, off by about a quarter of the true value.
In conclusion, while the model shows a good performance in some ways, especially when below 250 counts, there are limitations. It produced concentrated predictions that don’t reflect the true distribution’s variability and it was also focused heavily on certain values.
@@ V Conclusion
The success of DeepSeek R1 highlights the effectiveness of RL in enhancing the reasoning abilities of LLMs for complex tasks, such as solving math problems. This approach has been extended to VLMs for tasks like interpreting mathematical figures. Previous work often used a simplistic ”0/1” reward model, reducing answers to binary outcomes like ”YES/NO”. However, this reward function is inadequate for problems that don’t require exact matches, such as crowd counting or position matching with intersection over union (IOU).
To address this, we proposed FGRPR, which leverages the robust RL method, GRPO, combined with a fuzzy reward system that assigns higher rewards for values closer to the ground truth. Our experiments demonstrate that the traditional ”0/1” reward model with GRPO underperforms compared to SFT in these tasks. In contrast, FGRPR shows superior performance across all models, even surpassing GPT-4o and LLaMA-90B in in-domain datasets.
While FGRPR performs comparably to SFT in out-of-domain datasets with smaller target values, it excels with larger targets. This makes sense since traditional RL methods disregard the distinction between $\frac{y}{2}$ and $\frac{3y}{2}$. When $y$ is smaller, although our method remains competitive, it doesn’t always provide an advantage where can be improved in the future.

### v2

@@ IV-A Dataset
The dataset used in this article is derived from existing datasets for counting tasks. The training dataset includes aerial sheep [38], video game characters from GTAV_Head [39], street pedestrians [40], wheat heads from the Global Wheat Head Detection (GWHD) dataset [41], and cars from TRANCOS [42]. The testing dataset includes an additional dataset, manatee [43], for out-of-domain evaluation.
@@
Within our dataset, not all the images from the original datasets are used. If original dataset has train and test sets, then the original train and test samples are picked randomly to use as part of our own train and test sets, respectively. If the original dataset is not split into train and test sets, we first divide it into two parts, and then select images from these subsets to form our train and test sets. In our dataset, we use up to 1000 training images and 100 testing images.
Table I provides a summary of the datasets used in our study, highlighting the diversity and complexity of the counting task. The dataset encompasses a wide range of scenarios, from relatively simple images containing only a few objects to highly complex scenes with thousands of densely packed, overlapping instances. The variation in object counts, as indicated by the minimum, maximum, and mean values, demonstrates the challenge of accurately estimating quantities across different contexts. Some datasets, such as the video game characters and street pedestrians, include images where objects are densely clustered, making separation and identification difficult. Additionally, environmental noise, including buildings, trees, and other background elements, further complicates detection in real-world scenes. The dataset complexity indicates the challenges of performing a counting task with a single model.
@@ IV-B1 Models
  • Baseline Models: We evaluate several widely used benchmark models, including GPT-4o-2024-11-20 (GPT-4o), Claude Sonnet-3.5-2024-06-20 (Sonnet-3.5), LLaMA 2-70B (LLaMA-70B), Intern-VL-26B (In-VL2-26B), Qwen2.5-VL-3B (Q-3B), and Qwen2.5-VL-7B (Q-7B).
@@
  • GRPO Models: We conduct full-scale reinforcement learning training with the GRPO method on Qwen2.5-VL-3B (Q-3B-R1) and Qwen2.5-VL-7B (Q-7B-R1), employing the standard accuracy-based reward defined in equation 3.
  • CrowdVLM-R1 Models (Ours): We also perform full-scale RL training using our proposed FGRPR method on Qwen2.5-VL-3B (Q-3B-FR) and Qwen2.5-VL-7B (Q-7B-FR). Our fuzzy reward function is defined in equation 4.
@@ IV-B2 Experimental Environment
During training and testing for GRPO and FRGPR, we use 8 Nvidia L40S GPUs (each with 46GB of memory). The training dataset consists of 3803 images and 5 classes. The experiment is set to have a max prompt length of 1024, number of generations 8, per-device training batch size 6, gradient accumulation steps 2, data types bf16, and epochs 2. The testing consists of 600 images and 6 classes (one class is out of training samples). We use a batch size of 2 and data types bf16. It is worth noting that we use the same training and testing hyperparameters and prompt method across all the reinforcement learning models.
@@ IV-C Evaluation Metrics
To evaluate the performance of the crowd counting task, we employ Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), which are standard metrics in the field. MAE measures the average absolute difference between predicted and actual crowd counts, providing an intuitive measure of accuracy:
@@
To comprehensively assess model performance, we employ additional evaluation metrics, including the coefficient of determination ($R^{2}$), Mean Absolute Percentage Error (MAPE), Pearson’s correlation coefficient, and Spearman’s rank correlation coefficient. Their definitions are provided below.
IV-C1 Coefficient of Determination $R^{2}$
The $R^{2}$ score measures how well the predicted values fit the actual values, indicating the proportion of variance explained by the model:
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}}{\sum{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (10)
where $y_{i}$ and $\hat{y_{i}}$ denote the actual and predicted values, respectively, $\bar{y}$ is the mean of the actual values, and $n$ is the total number of samples. An $R^{2}$ value closer to 1 indicates a better fit.
IV-C2 Mean Absolute Percentage Error (MAPE)
MAPE quantifies the relative prediction error as a percentage, providing an intuitive measure of accuracy:
$\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100$ (11)
A lower MAPE value indicates higher prediction accuracy. However, it is sensitive to small actual values, which can lead to inflated error percentages.
IV-C3 Pearson’s Correlation Coefficient
Pearson’s correlation coefficient ($r$) measures the linear relationship between predicted and actual values, ranging from -1 to 1:
$r=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})(\hat{y_{i}}-\bar{\hat{y}})}{\sqrt{\sum{i=1}^{n}(y_{i}-\bar{y})^{2}}\sqrt{\sum_{i=1}^{n}(\hat{y}_{i}-\bar{\hat{y}})^{2}}}$ (12)
A value close to 1 indicates a strong positive correlation, while a value close to -1 suggests a strong negative correlation.
IV-C4 Spearman’s Rank Correlation Coefficient
The Spearman’s correlation ($\rho$) assesses the monotonic relationship between predicted and actual values by ranking the data points:
$\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)}$ (13)
where $d_{i}$ is the difference between the ranks of $y_{i}$ and $\hat{y_{i}}$, and $n$ is the number of observations. Unlike Pearson’s correlation, Spearman’s correlation captures non-linear relationships by evaluating rank similarity.
Figure 5: Training reward curves for different models show the same trend, while for precision reward, bigger models dominate the smaller models within the same method over the training steps after smoothed. And their format reward lines do not have any difference after 100 steps.
@@ IV-D Results
The overall model performance across different domains is shown in Table II. Green backgroundindicates the best results, while light greenrepresents the second-best results. Among the baseline models, LLaMA-70B demonstrates outstanding performance in four of the six domain datasets. Nevertheless, GPT-4o and Sonnet-3.5 achieve lower MAE and RMSE scores compared to LLaMA-70B. Specifically, GPT-4o exhibits comparable performance to LLaMA-70B on the out-of-domain dataset (Manatees), displaying an MAE of approximately 5.7 and an RMSE around 10. On the other hand, Q-3B and Q-7B consistently show among the poorest metric results across all six datasets.
@@
However, after traditional supervised fine-tuning, both Q-3B-SFT and Q-7B-SFT improve significantly. Specifically, they outperform not only their base versions but also all baseline models, including state-of-the-art models such as GPT-4o and LLaMA-70B, across five in-domain datasets. Although Q-7B-SFT achieves a lower overall MAE compared to Q-3B-SFT, its performance is not consistently better across these five in-domain datasets. Furthermore, Q-7B-SFT presents a higher RMSE value, indicating that its prediction errors exhibit greater variance and instability.
@@
After fine-tuning with GRPO methods employing the traditional $0/1$ reward function (rewarding predictions with $1$ if $\frac{\hat{y}-y}{y}<0.5$, otherwise 0), the results indicate that although this approach exhibits some effectiveness, the models Q-3B-R1 and Q-7B-R1 show worse performance than their corresponding directly fine-tuned counterparts, Q-3B-SFT and Q-7B-SFT. This outcome reveals a limitation of applying such a GRPO-based reward approach specifically for the crowd counting task.
@@
Models trained with our proposed FGRPR method, particularly Q-7B-FR, exhibit state-of-the-art performance across all five in-domain datasets. They continue to reduce the MAE and RMSE compared to the previously best-trained model, Q-7B-SFT, showing consistency in error reduction. For the 3B variant, Q-3B-FR achieves better performance than the previous SFT model (Q-3B-SFT) in two datasets, slightly worse performance in sheep counting, and equivalent performance in wheat head and traffic car counting datasets. Moreover, for the out-of-domain dataset, SFT and FGRPR show similar performance, which improved compared to their base models that show better performance than bigger models In-VL2-26B and Sonnet-3.5, but still can not achieve the best industry models like GPT-4o and LLaMA-70B. Considering the small size of our base models (only 3B and 7B), these results collectively prove the effectiveness of our proposed FGRPR reward function.
@@
This effectiveness is further supported by Table III, where the models are evaluated based on different count ranges. In in-domain datasets, the SFT and FGRPR methods have similar performance with fewer objects, in ranges [1-20] and [21-50]. And all beat the most recent vision language and open-vocabulary counting method, T-Rex[44], in out of domain dataset. However, as object counts exceed 50, FGRPR shows superior performance compared to SFT and GRPO, which is guided by the accuracy reward, equation 3. This highlights the effectiveness of our proposed reward function, equation 7, particularly the component ($\frac{\left|\hat{y}-y\right|}{y}$), which offers higher rewards for more accurate counts.
@@ IV-E1 Model rewards
In Fig. 5, all four models demonstrate an upward trend in both precision and format rewards, as shown in the first-row subfigures where a simple moving average with a window size of 20 is applied across 600 training steps. The 7B models start at higher points than the 3B models, indicating superior initial performance, consistent with prior studies that larger models typically outperform smaller ones within the same family. And 3B model with FGRPR has similar start values to Q-7B-R1.
All four models display a sharp increase within the initial 100 training steps, followed by fluctuating yet steady improvement in precision rewards over the subsequent 500 steps, reaching their peak around step 600. Notably, the FGRPR method demonstrates a sharper and steadier growth trajectory compared to the traditional GRPO method. Furthermore, the continuous upward trend shown by the FGRPR model at the training endpoint suggests the potential for further improvement with additional training data or a longer training period, whereas models trained via GRPO with traditional rewards plateau and exhibit little to no increase in reward values during the second half of training. Meanwhile, all four models have identical format reward curves, as they share the same format reward function.
@@ IV-E2 Concentrated results
Experiment results show that when the number of detected objects within the images is under 250, models prefer to predict some fixed count values.
@@
Fig. 6 presents distribution statistics for ground truth over prediction results. The histogram and density plots show that both distributions are highly right-skewed, with the majority of values concentrated below 250. However, the second row subfigures dramatically illustrate disparity between GT and predictions - GT values show numerous outliers spanning a much wider range, while predicted values display fewer outliers and compressed variability.
@@
Table IV reveals additional metrics for the prediction of the model. The $R^{2}$ value of 0.8795 indicates that the model explains nearly 88% of the variance. Strong correlation coefficients (Pearson = 0.9655, Spearman = 0.9497) confirm that the model generally preserves rank ordering. And an MAPE of 23.89% suggests predictions are, on average, off by about a quarter of the true value.
In conclusion, while the model shows a good performance in some ways, especially when below 250 counts, there are limitations. It produces concentrated predictions that do not reflect the true distribution’s variability, and it also focuses heavily on certain values.
@@ V Conclusion
The success of DeepSeek R1 highlights the effectiveness of RL in enhancing the reasoning abilities of LLMs for complex tasks, such as solving math problems. This approach has been extended to VLMs for tasks like interpreting mathematical figures. Previous work often used a simplistic ”0/1” reward model, reducing answers to binary outcomes like ”YES/NO”. However, this reward function is inadequate for problems that require exact matches, such as crowd counting or position matching with intersection over union.
To address this, we proposed FGRPR, which leverages the robust RL method, GRPO, combined with a fuzzy reward system that assigns higher rewards for values closer to the ground truth. Our experiments demonstrate that the traditional ”0/1” reward model with GRPO underperforms compared to SFT in these tasks. In contrast, FGRPR shows superior performance across all models, even surpassing GPT-4o and LLaMA-70B in in-domain datasets.
While FGRPR performs comparably to SFT in out-of-domain datasets with smaller target values, it excels with larger targets. This makes sense since traditional RL methods disregard the distinction between $\frac{y}{2}$ and $\frac{3y}{2}$. When $y$ is smaller, although our method remains competitive, it does not always provide an advantage—an aspect that could be improved in the future.
@@
Acknowledgment
This work has been supported in part by the US National Science Foundation (NSF) under Grant No. IOS-2430224.